OpenStudy (anonymous):

Let r = 1 + cos θ. Find the horizontal and vertical tangent lines to the polar curve.

OpenStudy (anonymous):

here, like usual, we are trying to find points where \(\Large {dy\over dx}=0 \text{ or }\infty\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for horizontal tangent line it'd just be dy= 0 which i'm getting values for theta to be 0, 2pi, pi/2, and 3pi/s

OpenStudy (anonymous):

using chain rule, we have: \[ \Large{dy\over dx}=\frac{{dr\over d\theta}\sin\theta+r\cos\theta}{{dr\over d\theta}\cos\theta-r\sin\theta}\\ {dy\over dx}=\frac{(-\sin\theta)\sin\theta+(1+\cos\theta)\cos\theta}{(-\sin\theta)\cos\theta-(1+\cos\theta)\sin\theta} \] use this now, to find for what \(\theta\) is the derivative "0" or infinite

OpenStudy (anonymous):

HORIZONTAL tangent: \[-\sin^2\theta+\cos\theta+\cos^2\theta=0\\ (1-\cos^2\theta)+\cos\theta+\cos^2\theta=0\\ 2\cos^2\theta+\cos\theta-1=0\] solve this and find the possilbe "theta"

OpenStudy (anonymous):

then you can get the equation of horizontal tangent.

OpenStudy (anonymous):

similarly, for the vertical, set the denominator to zero ad solve

OpenStudy (anonymous):

for dx i'm getting -sin(theta) -2cos(theta)sin(theta)

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

what do you mean by "for dx"?